OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
BHS4
Scheduled Maintenance Report for Network & Infrastructure
Completed
Because of an electrical issue we have lost network devices.


Update(s):

Date: 2017-10-23 01:01:17 UTC
Lasts servers will restart soon.

Date: 2017-10-23 00:49:20 UTC
New eletrical system has been installed.
Last switches are coming back online

Date: 2017-10-23 00:37:26 UTC
Access switches are still down and Electrical repair are still ongoing.

Date: 2017-10-22 23:36:56 UTC
bhs4-14ab-n56 is OK now

Date: 2017-10-22 23:09:11 UTC
Technicians need to bring a second power source.
This can take 1 hour

Date: 2017-10-22 22:48:43 UTC
All PCI is OK
bhs-z2b1-a70 OK
bhs-z2b2-a70 OK
bhs-z2b3-a70 OK
bhs-z2b4-a70 OK
bhs-z2b5-a70 OK
bhs-z2b6-a70 OK
bhs-z2b10-a70 OK
bhs-z2b11-a70 OK
bhs-z2b12-a70 OK
bhs-z2b13-a70 OK
bhs-z2b14-a70 OK
bhs-z2b15-a70 OK
bhs-z2b18-a70 OK
bhs-z2c1-a70 OK
bhs-z2c2-a70 OK
bhs-z2c3-a70 OK
bhs-z2s5-a70 OK


bhs4-11a-n56 OK
bhs4-14a-n56



bhs4-100-n56-vrack
fex150
fex157
fex158
fex159

Date: 2017-10-22 22:25:42 UTC
bhs-z2b1-a70 OK
bhs-z2b2-a70 OK
bhs-z2b3-a70 OK
bhs-z2b4-a70
bhs-z2b5-a70
bhs-z2b6-a70
bhs-z2b10-a70 OK
bhs-z2b11-a70
bhs-z2b12-a70 OK
bhs-z2b13-a70 OK
bhs-z2b14-a70 OK
bhs-z2b15-a70 OK
bhs-z2b18-a70
bhs-z2c1-a70 OK
bhs-z2c2-a70 OK
bhs-z2c3-a70 OK
bhs-z2s5-a70


bhs4-11a-n56 OK
bhs4-14a-n56



bhs4-100-n56-vrack
fex150
fex157
fex158
fex159

Date: 2017-10-22 21:57:42 UTC
We will plug switches to a secondary power source

Date: 2017-10-22 21:53:19 UTC
Some Vrack FEX are also impacted

Date: 2017-10-22 21:48:38 UTC
SD impacted switches:
bhs4-11ab-n56 bhs4-14ab-n56

Date: 2017-10-22 21:48:09 UTC
PCI VPS impacted switches:
bhs-z2b1-a70
bhs-z2b2-a70
bhs-z2b3-a70
bhs-z2b4-a70
bhs-z2b5-a70
bhs-z2b6-a70
bhs-z2b10-a70
bhs-z2b11-a70
bhs-z2b12-a70
bhs-z2b13-a70
bhs-z2b14-a70
bhs-z2b15-a70
bhs-z2b18-a70

bhs-z2c1-a70
bhs-z2c2-a70
bhs-z2c3-a70

bhs-z2s5-a70
Posted Oct 22, 2017 - 21:47 UTC